### v8

@@ 4.6 Generalization Experiments for the Proposed Methods
Model Param(M) GFLOPs mAP50 mAP50:95 APs APm
@@ 5 Conclusions
In this paper, we proposed TinyDef-DETR, a DETR-based framework tailored for defect detection in UAV imagery of transmission lines. By integrating an Edge-Enhanced ResNet backbone, a stride-free space-to-depth module, a cross-stage dual-domain multi-scale attention mechanism, and a Focaler-Wise-SIoU regression loss, the model effectively addresses the challenges posed by small, ambiguous, and background-cluttered defects.
Extensive experiments on the CSG-ADCD dataset, as well as real-world UAV imagery, demonstrated that TinyDef-DETR consistently outperforms representative detectors in terms of detection accuracy, recall, and robustness, while maintaining computational efficiency. Ablation studies further verified the effectiveness of each proposed component, highlighting their complementary contributions to the overall performance.
The results confirm that TinyDef-DETR not only achieves strong detection capability for small and difficult objects but also provides a practical and efficient solution for real-world UAV-based inspection scenarios. Given its modular and lightweight design, the proposed framework holds promise for broader applications in power system inspection and can be extended to other small-object detection tasks in aerial remote sensing.

### v9

@@ 4.7 Generalization Experiments for the Proposed Methods
Model Params (M) GFLOPs mAP50 mAP50:95 APs APm
@@
5 Discussion
5.1 Theoretical Mechanisms of Performance Enhancement
The performance gains observed with TinyDef-DETR arise from the mutually reinforcing effects of its architectural and loss-design innovations, each addressing a specific bottleneck in UAV-based defect detection. At the representational level, the edge-enhanced ResNet backbone strengthens the model’s sensitivity to fine structural cues by amplifying high-frequency components associated with defect boundaries. This modification counteracts the intrinsic attenuation of edge information caused by deep convolutional downsampling, enabling the model to preserve discriminative contours essential for identifying small or visually ambiguous defects. From a theoretical standpoint, such enhancement effectively increases the signal-to-noise ratio for fine-grained features and improves the separability of defect patterns within the feature manifold.
The stride-free space-to-depth (SPD) module introduces a second source of improvement by performing spatial reduction without discarding pixel-level detail. Unlike conventional stride-based downsampling, which induces aliasing and information loss, SPD reorganizes high-resolution inputs into channel-enriched tensors. This preserves local texture, minimizes feature degradation for small targets, and stabilizes gradient flow. Theoretically, this operation enhances the mutual information between shallow and deep features, enabling a more coherent propagation of small-object descriptors throughout the network hierarchy.
Furthermore, the cross-stage dual-domain multi-scale attention mechanism integrates spatial and frequency-domain representations to model global context and local variations more holistically. By enabling multi-scale interactions across network stages, this module addresses the scale variability characteristic of UAV imagery and improves the model’s ability to capture long-range structural dependencies. In essence, the mechanism expands the effective receptive field while simultaneously preserving the locality required for small-object detection, creating a balanced and robust feature fusion process.
Finally, the Focaler–Wise–SIoU loss contributes an optimization-level enhancement by modulating regression difficulty through an EMA-normalized difficulty indicator. The unimodal weighting function prioritizes moderately hard samples while attenuating extreme outliers and trivial cases, resulting in a smoother and more stable error landscape. This adaptive reweighting improves convergence behavior, enhances bounding-box localization for small defects, and mitigates gradient domination by noisy samples. From a theoretical perspective, the loss introduces a dynamic sample-importance prior that aligns training emphasis with statistically informative examples, thereby improving both robustness and generalization.
Collectively, these mechanisms form a coherent theoretical foundation explaining why TinyDef-DETR exhibits strong performance on small, ambiguous, and cluttered defect instances in UAV imagery. Each component contributes a targeted enhancement—representational fidelity, information preservation, multi-scale context integration, and difficulty-aware optimization—that together yield a substantial improvement over conventional detection pipelines.
5.2 Limitations and Future Directions
Despite its effectiveness, TinyDef-DETR still exhibits certain limitations that warrant further exploration. First, although SPD alleviates information loss during downsampling, the increased channel dimensionality demands additional memory and may introduce latency on resource-limited embedded UAV platforms. Lightweight kernel decomposition or dynamic channel pruning may help reduce overhead while preserving detail fidelity.
Second, the dual-domain attention module, while beneficial for multi-scale fusion, introduces additional complexity to the training pipeline. In scenarios involving extremely high-resolution imagery or large-scale datasets, the global context modeling may become computationally expensive. Future work may explore hierarchical or sparse attention variants to reduce computational burden while maintaining cross-stage consistency.
Third, the difficulty-modulated regression loss relies on reliable EMA statistics of IoU distributions. In datasets with extreme imbalance or rapidly shifting difficulty patterns, the moving average may adapt too slowly, reducing the responsiveness of the modulation mechanism. Developing more flexible or instance-adaptive normalization strategies could further stabilize optimization and improve performance on highly heterogeneous data.
Additionally, although the present study focuses on transmission line defect detection, the generalization of TinyDef-DETR to other aerial inspection domains—such as railway, bridge, and pipeline monitoring—remains to be systematically evaluated. Cross-domain validation, domain adaptation methods, or self-supervised pretraining may help broaden its applicability.
Finally, the current framework does not explicitly incorporate geometric priors, structural constraints, or transformer-based depth modeling that could further improve detection under severe occlusion or low-contrast imaging. Integrating 3D contextual cues, physical-structure modeling, or multi-modal inputs (e.g., infrared or LiDAR) could provide additional robustness in complex operational environments.
@@ 6 Conclusions
In this paper, we introduced TinyDef-DETR, a DETR-based framework tailored to defect detection in UAV imagery of transmission lines. The design combines an Edge-Enhanced ResNet backbone, a stride-free space-to-depth (SPD) downsampling module, a cross-stage dual-domain multi-scale attention mechanism, and a Focaler–Wise–SIoU regression loss. Together, these components target the core difficulties of this task—tiny object scale, visual ambiguity, and background clutter—by strengthening edge/detail representation, preserving pixel-level cues during downsampling, enhancing global–local feature interaction, and stabilizing box regression for small targets.
Comprehensive experiments on the CSG-ADCD dataset and real-world UAV imagery show that TinyDef-DETR consistently surpasses representative detectors in detection accuracy, recall, and robustness, while maintaining competitive computational efficiency. Ablation studies further confirm that each module contributes complementary gains to the final system, clarifying the role of edge-aware features, lossless downsampling, dual-domain attention, and shape/angle-aware localization in the overall performance.
These results indicate that TinyDef-DETR delivers strong detection capability for small and challenging defects and is well-suited for practical UAV-based inspection. Owing to its modular and lightweight design, the framework is readily adaptable to broader power-system inspection workflows and can be extended to other small-object detection problems in aerial and remote-sensing applications.